# Aperiodic sets of three types of convex polygons

### v1

@@ Appendix D.4 Conditions for treating anterior and posterior sides as the same type
The ASP created by $S_{t}(\alpha,\beta)$ that satisfies the restriction on the number of types of congruent diagrams can be regarded as “an ASP comprising three types of convex polygons” from the viewpoint that does not distinguish between anterior and posterior sides.
If $S_{t}(\alpha,\beta)$ is subject to the same rules as the monohedral tiling, that is, “the use of reflected tiles is allowed in the tiling,” then $S_{t}(\alpha,\beta)$ is not ASP. Tile(1,1)11(1,1)( 1 , 1 ) can generate a periodic tiling if the use of reflected tiles is allowed during the tiling generation process [9]. In other words, if the reflected $CP(X_{14})$ is allowed in tiling, the set comprises of six types of convex polygons that can generate periodic tiling.
On the contrary, $S_{t}(\alpha,\beta)$ considered the following “tile combination specification condition” as a method to be able to argue that “ASP comprises of three types of convex polygons” from the viewpoint that does not distinguish between anterior and posterior sides.
@@
Moreover, “convex polygons that cannot be reflected during the tiling generation process are assumed not to have the property of line symmetry” is important. As mentioned previously, the set of six convex polygons with $S_{t}(\alpha,\beta)$ and reflected $CP(X_{14})$ can form periodic tiling; however, it may also be possible to form periodic tiling in other cases. For example, for the case of $S_{t}(\alpha=105^{\circ},\beta=75^{\circ})$ in Figure 29, if $CP(X_{14})$, $CP(X_{3})$, and reflected $CP(X_{14})$ are used, periodic tiling is possible, as shown in Figure 32. In this case, the merged diagram of $CP(X_{14})$ and reflected $CP(X_{14})$ can be regarded as forming a convex octagon. In other words, Figure 32 can be regarded as periodic tiling using anterior and posterior convex octagons and convex pentagons without any distinction between the anterior and posterior sides (two types of convex pentagons and octagons that are not monotiles from a viewpoint that does not distinguish between the anterior and posterior sides).
@@ Appendix D.5 Supplement 1
In addition to Figure 26, there is another method for dividing Tile(1,1)11(1,1)( 1 , 1 ) such that the number of convex polygonal types is less than or equal to three from a viewpoint that does not distinguish between the anterior and posterior sides, as shown in Figure 33. However, the sets of convex polygons created by the divisions in Figure 33 always contain convex pentagonal monotiles. The convex pentagon monotile belonging to both the Type 2 and Type 4 families with line symmetry can form periodic tiling (see Appendix A). For the division that makes the convex pentagonal monotiles belonging to the Type 13 family in Figure 33, the tile set of the three types of convex polygons contains the anterior and posterior sides of convex pentagonal monotiles belonging to the Type 13 family. Therefore, the tile sets of convex polygons shown in Figure 33 are not ASP.
@@ Appendix D.6 Supplement 2
As shown in Figure 36, we found that there is another method for dividing Tile(1,1)11(1,1)( 1 , 1 ) such that the number of convex polygonal types is less than or equal to three from a viewpoint that does not distinguish between the anterior and posterior sides. In Figure 36, the point $Q$ moves on the edge $X_{5}X_{6}$ and the point $T$ moves on the edge $X_{12}X_{13}$ of Tile(1,1)11(1,1)( 1 , 1 ). In Figure 36, the convex pentagon $CP_{2}(X_{3})$ with the points “$X_{3}$, $X_{4}$, $X_{5}$, $X_{1}$, $X_{2}$” as vertices is fixed shape. Additionally, the convex pentagon $CP_{2}(X_{7})$ with the points “$X_{7}$, $X_{8}$, $R$, $Q$, $X_{6}$” as vertices, the convex pentagon $CP_{2}(X_{9})$ with the points “$X_{9}$, $X_{10}$, $S$, $R$, $X_{8}$” as vertices, and the convex pentagon $CP_{2}(X_{11})$ with the points “$X_{11}$, $X_{12}$, $T$, $S$, $X_{10}$” as vertices are congruent convex pentagons. However, $CP_{2}(X_{7})$ and $CP_{2}(X_{11})$ have a reflection relation with $CP_{2}(X_{9})$. The convex polygon $CP_{2}(X_{14})$ with the points “$X_{14}$, $X_{1}$, $X_{5}$, $Q$, $R$, $S$, $T$” as vertices is a heptagon, but it is a hexagon if $\gamma=90^{\circ}$ and $\delta=120^{\circ}$.
@@
      Case 5. When $\gamma=90^{\circ}$ and $\delta=120^{\circ}$ in Figures 36, there is a “set of three types of convex pentagons ($CP_{2}(X_{3})$, $CP_{2}(X_{7})$, $CP_{2}(X_{9})$, $CP_{2}(X_{11})$) and one type of convex hexagon ($CP_{2}(X_{14})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is a “set of two types of convex pentagons and one type of convex hexagon.”
@@
      Case 6. When $\gamma\neq 90^{\circ}$ and $\delta\neq 120^{\circ}$ in Figures 36, there is a “set of three types of convex pentagons ($CP_{2}(X_{3})$, $CP_{2}(X_{7})$, $CP_{2}(X_{9})$, $CP_{2}(X_{11})$) and one type of convex heptagon ($CP_{2}(X_{14})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is a “set of two types of convex pentagons and one type of convex heptagon.”

### v3

@@ Appendix D.4 Conditions for treating anterior and posterior sides as the same type
The ASP created by $S_{t}(\delta,\epsilon)$ that satisfies the restriction on the number of types of congruent diagrams can be regarded as “an ASP comprising three types of convex polygons” from the viewpoint that does not distinguish between anterior and posterior sides.
If $S_{t}(\delta,\epsilon)$ is subject to the same rules as the monohedral tiling, that is, “the use of reflected tiles is allowed in the tiling,” then $S_{t}(\delta,\epsilon)$ is not ASP. Tile(1,1)11(1,1)( 1 , 1 ) can generate a periodic tiling if the use of reflected tiles is allowed during the tiling generation process [9]. In other words, if the reflected $CP(X_{14})$ is allowed in tiling, the set comprises of six types of convex polygons that can generate periodic tiling.
On the contrary, $S_{t}(\delta,\epsilon)$ considered the following “tile combination specification condition” as a method to be able to argue that “ASP comprises of three types of convex polygons” from the viewpoint that does not distinguish between anterior and posterior sides.
@@
Moreover, “convex polygons that cannot be reflected during the tiling generation process are assumed not to have the property of line symmetry” is important. As mentioned previously, the set of six convex polygons with $S_{t}(\delta,\epsilon)$ and reflected $CP(X_{14})$ can form periodic tiling; however, it may also be possible to form periodic tiling in other cases. For example, for the case of $S_{t}(\delta=105^{\circ},\epsilon=75^{\circ})$ in Figure 32, if $CP(X_{14})$, $CP(X_{3})$, and reflected $CP(X_{14})$ are used, periodic tiling is possible, as shown in Figure 35. In this case, the merged diagram of $CP(X_{14})$ and reflected $CP(X_{14})$ can be regarded as forming a convex octagon. In other words, Figure 35 can be regarded as periodic tiling using anterior and posterior convex octagons and convex pentagons without any distinction between the anterior and posterior sides (two types of convex pentagons and octagons that are not monotiles from a viewpoint that does not distinguish between the anterior and posterior sides).
@@ Appendix D.5 Supplement 1
In addition to Figure 29, there is another method for dividing Tile(1,1)11(1,1)( 1 , 1 ) such that the number of convex polygonal types is less than or equal to three from a viewpoint that does not distinguish between the anterior and posterior sides, as shown in Figure 36. However, the sets of convex polygons created by the divisions in Figure 36 always contain convex pentagonal monotiles. The convex pentagonal monotile belonging to both the Type 2 and Type 4 families with line symmetry can form periodic tiling (see Appendix A). For the division that makes the convex pentagonal monotiles belonging to the Type 13 family in Figure 36, the tile set of the four types of convex polygons contains the anterior and posterior sides of convex pentagonal monotiles belonging to the Type 13 family. Therefore, the tile sets of convex polygons shown in Figure 36 are not ASP.
@@ Appendix D.6 Supplement 2
As shown in Figure 39, we found that there is another method for dividing Tile(1,1)11(1,1)( 1 , 1 ) such that the number of convex polygonal types is less than or equal to three from a viewpoint that does not distinguish between the anterior and posterior sides. In Figure 39, the point $Q$ moves on the edge $X_{5}X_{6}$ and the point $T$ moves on the edge $X_{12}X_{13}$ of Tile(1,1)11(1,1)( 1 , 1 ). In Figure 39, the convex pentagon $CP_{2}(X_{3})$ with the points “$X_{3}$, $X_{4}$, $X_{5}$, $X_{1}$, $X_{2}$” as vertices is fixed shape. Additionally, the convex pentagon $CP_{2}(X_{7})$ with the points “$X_{7}$, $X_{8}$, $R$, $Q$, $X_{6}$” as vertices, the convex pentagon $CP_{2}(X_{9})$ with the points “$X_{9}$, $X_{10}$, $S$, $R$, $X_{8}$” as vertices, and the convex pentagon $CP_{2}(X_{11})$ with the points “$X_{11}$, $X_{12}$, $T$, $S$, $X_{10}$” as vertices are congruent convex pentagons. However, $CP_{2}(X_{7})$ and $CP_{2}(X_{11})$ have a reflection relation with $CP_{2}(X_{9})$. The convex polygon $CP_{2}(X_{14})$ with the points “$X_{14}$, $X_{1}$, $X_{5}$, $Q$, $R$, $S$, $T$” as vertices is a heptagon, but it is a hexagon if $\zeta=120^{\circ}$ and $\eta=90^{\circ}$.
@@
      Case 5. When $\zeta=120^{\circ}$ and $\eta=90^{\circ}$ in Figures 39, there is a “set of three types of convex pentagons ($CP_{2}(X_{3})$, $CP_{2}(X_{7})$, $CP_{2}(X_{9})$, $CP_{2}(X_{11})$) and one type of convex hexagon ($CP_{2}(X_{14})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is a “set of two types of convex pentagons and one type of convex hexagon.”
@@
      Case 6. When $\zeta\neq 120^{\circ}$ and $\eta\neq 90^{\circ}$ in Figures 39, there is a “set of three types of convex pentagons ($CP_{2}(X_{3})$, $CP_{2}(X_{7})$, $CP_{2}(X_{9})$, $CP_{2}(X_{11})$) and one type of convex heptagon ($CP_{2}(X_{14})$),” and, from the viewpoint that does not distinguish between anterior and posterior sides, it is a “set of two types of convex pentagons and one type of convex heptagon.”